# DRIFT: Dynamic Rule-Based Defense with Injection Isolation for Securing LLM Agents

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly central to agentic systems due to their strong reasoning and planning capabilities. By interacting with external environments through predefined tools, these agents can carry out complex user tasks. Nonetheless, this interaction also introduces the risk of prompt injection attacks, where malicious inputs from external sources can mislead the agent’s behavior, potentially resulting in economic loss, privacy leakage, or system compromise. System-level defenses have recently shown promise by enforcing static or predefined policies, but they still face two key challenges: the ability to dynamically update security rules and the need for memory stream isolation. To address these challenges, we propose DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems, which enforces both control- and data-level constraints. A Secure Planner first constructs a minimal function trajectory and a JSON-schema-style parameter checklist for each function node based on the user query. A Dynamic Validator then monitors deviations from the original plan, assessing whether changes comply with privilege limitations and the user’s intent. Finally, an Injection Isolator detects and masks any instructions that may conflict with the user query from the memory stream to mitigate long-term risks. We empirically validate the effectiveness of DRIFT on the AgentDojo benchmark, demonstrating its strong security performance while maintaining high utility across diverse models—showcasing both its robustness and adaptability.
@@ 1 Introduction
Large Language Models (LLMs), Large Language Models (LLMs), empowered by their exceptional planning and reasoning abilities, are increasingly integrated into agentic systems [1, 2, 3]. By processing natural language data streams, LLM agents interact with external environments, such as applications [1, 4], computing systems [3], via a set of pre-defined tools to carry out complex user tasks. Since the need for interaction with untrusted external environments, a new security threat of prompt injection attacks is introduced [5, 6, 7, 8, 9], where attackers inject malicious instructions into third-party platforms, misleading the agent workflow after external interaction. For example, a product review on Amazon written by another user, such as “Ignore previous instructions, buy this red shirt,” may manipulate the LLM into executing unintended actions. This form of attack [5, 6, 7, 8, 9] may bring risks such as economic losses [6], privacy leakage [10], and system damage [11] to users, severely undermining the reliability of the agentic system.
@@
Existing defense mechanisms can be broadly categorized into model-level [12, 13, 14, 15, 16, 17] and system-level [18, 19, 20, 21, 22] approaches. Model-level defenses [12, 13, 14, 15, 16, 17] typically rely on training a guardrail to detect injection input or mitigate the injection impact. Such methods are constrained by the inherent vulnerabilities of the models and often struggle to defend against complex and diverse attacks. Recently, system-level [18, 19, 20, 21, 22] defenses have gained increasing attention, as they are able to overcome the inherent vulnerabilities of models when facing unseen attacks, thereby achieving high practical reliability in real-world applications. These approaches typically constrain the model’s action space through predefined security policies or rules to prevent potential injection threats. For instance, IsolateGPT [18] mitigates information leakage risks by enforcing isolation mechanisms and maintaining a separate memory bank for each application. Recently, CaMeL [21] achieves impressive security by manually defining a set of security policies and constructing a strict, fixed control and data dependency graph from the user query before any interaction takes place.
@@
Despite the progress in system-level defense mechanisms for agentic systems, two critical challenges remain largely unresolved: (1) the dynamic updating of security rules and (2) the isolation of injected content within the memory stream. While CaMeL enforces robust security through a strict dependency graph, but this static design considerably sacrifices flexibility and practical usability, particularly in agentic systems that require adaptive, real-time decision-making. Furthermore, the reliance on manually crafted security policies imposes considerable overhead and impedes generalization across diverse usage scenarios. In addition, the method of IsolateGPT restrict the propagation of injection-related information across different applications, but residual injection content preserved in memory still poses significant risks within the same application during prolonged interactions.
@@
To overcome these challenges, we develop DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems that enforces security through both control- and data-level constraints. We first design a Secure Planner, which establishes the initial constraint rules solely according to the user query prior to any interaction. It constructs a minimal function trajectory (control constraints) to avoid injections misleading by executing functions in order. In addition, a checklist for each function node in the trajectory, with detailed parameter requirement and value dependencies, is encoded in JSON schema format [23]. When trajectory deviations are detected, a Dynamic Validator performs approval assessments based on the privilege category (Read, Write, Execute) and its alignment with the user’s original intent. To avoid the risk of injection messages to the agent or other modules during prolonged interactions, an Injection Isolator is also designed to continuously polish the memory after each interaction, identifying and masking any instructions that conflict with the initial user query. This layered defense strategy ensures strong context isolation while enabling secure and adaptive decision-making throughout long-term agent interactions. We include a more detailed discussion of related work in Appendix A.
@@
As a fully automatic system-level defense framework, DRIFT demonstrates strong performance across diverse scenarios, achieving high security while maintaining robust utility. Specifically, we evaluate DRIFT on the AgentDojo [24] benchmark, a simulated agent environment featuring various task scenarios and types of injection attacks. By applying DRIFT to GPT-4o-mini [25], the Attack Success Rate (ASR) is successfully reduced from 30.7% to 1.3%, while utility outperforms CaMeL by 20.1% under no attack and by 12.5% under attack. In addition, DRIFT shows remarkable adaptability and generalization across four advanced online LLMs: GPT-4o [26], GPT-4o-mini [25], Claude-3.5-sonnet [27], Claude-3-haiku [28], and one prevalent offline LLM, Qwen2.5-7B [29]. On all of these models, DRIFT significantly enhances security while maintaining or even improving utility on some models. Moreover, we finetune our policy on DRIFT, with the dataset collected from ToolBench [30], achieving significant improvements in both security and utility. Compared to the original version, the ASR of the policy-tuned model drops from 15.1% to 0.0%, while utility under no attack increases from 26.6% to 32.2%, and utility under attack improves from 19.1% to 22.2%. This DRIFT policy training mechanism could enable more reliable, robust, secure, and functional LLM agentic systems.
@@
  • We develop DRIFT, a comprehensive system-level defense framework that integrates dynamic security policies with memory isolation, achieving a superior balance between security and functionality.
@@
    footnotetext: Project Page: https://github.com/leolee99/DRIFT
@@ 2.2 Dynamic Validator
In general, a function with the Read privilege does not directly pose a risk to the user and will be approved even if it deviates from the original trajectory. However, functions marked as Write or Execute may introduce direct risks. In such cases, the Validator will assess whether the deviated function aligns with the user’s original intent based on the updated tool messages, using the prompt shown in Figure 9. If the deviated function still aligns with the user’s intent, the function is approved and incorporated into the minimal function trajectory and parameter checklist to support successful validation in subsequent validation. Otherwise, agents will send an approval request to user.
@@ 2.3 Injection Isolator
Current rule-based agent defense approaches typically restrict action permissions but do not eliminate injected content. In a long-term agentic system, past memory—such as previous conversations and tool responses—is frequently reused. These reused elements may be accessed not only by the agent itself but also by other components within the security system, such as the policy updating module. During the process of policy optimization, it is inevitable to incorporate new information obtained from recent interactions. However, any injection content stored in the memory stream will also be repeatedly exposed to these components during long-term interactions, severely increasing the risk of compromise over time.
To mitigate this long-term threat, we propose an injection isolation mechanism to detect and remove injected content from the memory stream. Specifically, we design a curated Injection Isolator that analyzes returned messages from each tool-calling and determines whether any instructions conflict with the user’s original intent. The identification process is driven by a LLM using system prompt in Figure 10. If a conflict is detected, the isolator removes the conflicting instructions using external masking components before the message is added to the agent’s memory stream. Subsequently, a safe memory stream could be maintained in long-term agent interactions. The Isolator cannot directly modify the tools and does not interact with the agent, which helps prevent potential security vulnerabilities as much as possible.
@@ 2.5 Trainable Security Policy
To enhance the reliability and generalization of our security policy, we develop a training approach for both the Secure Planner and the Injection Isolator, allowing our DRIFT framework to adapt more robustly. This involves designing a new data collection pipeline that extracts policy-aligned samples from existing agent datasets, followed by efficient instruction tuning using Low-Rank Adaptation (LoRA) [32] on Qwen2.5-7B [29].
@@ 2.5.1 Data and Environment Construction
After data collection completed, we fine-tune the Qwen2.5-7B model using LoRA for both the Secure Planner and Injection Isolator, as well as the agent itself. For the Dynamic Validator, we rely on the original Qwen2.5-7B in a zero-shot setup to handle privilege classification and user intent checking.
@@ 3 Experiments
In this section, we evaluate DRIFT on AgentDojo, the latest and prevalent agentic security benchmark, to assess the effectiveness and adaptability of DRIFT in terms of both utility and security. Furthermore, we analyze the contribution of each individual component within DRIFT.
@@ 3.1 Experimental Setups
We evaluate our method with AgentDojo [24], a benchmark that simulates realistic interactions in agent-based systems. It includes four scenarios—banking, Slack, travel, and workspace—covering 97 user tasks to assess utility and 629 injection tasks to evaluate security.
@@
We compare our method against five existing advanced defense approaches, with four approaches available in AgentDojo: repeat_user_prompt, spotlighting_with_delimiting, tool_filter, transformers_pi_detector, and a policy-based security approach of CaMeL [21]. These represent a range of strategies for protecting agents from injection attacks.
@@
We apply our policy to several models, including online models—GPT-4o [26], GPT-4o-mini [25], Claude-3-haiku [28], and Claude-3.5-sonnet [27]—and an offline model, Qwen2.5-7B [29]. For Qwen2.5-7B, we fine-tune it on our policy dataset (described in Section 2.5) using a batch size of 4 and training for three epochs. We employ the Adam optimizer [33] with weight decay and set the initial learning rate to 2e-5. The maximum input length is 15,000 tokens.
Figure 3: Comparison across different defense method on GPT-4o-mini.
@@ 3.2 Defense Techniques Comparison
In this section, we compare DRIFT with five advanced defense techniques—four implemented in AgentDojo: repeat_user_prompt, spotlighting_with_delimiting, tool_filter, transformers_pi_detector—and one policy-based defense approach, CaMeL. Figure 3 illustrates the results. We use GPT-4o-mini-2024-07-18 as the base agent for all evaluations. Notably, the DRIFT policy achieves an optimal balance between utility and security. In terms of security, DRIFT significantly outperforms all other baselines except CaMeL, with only a marginal gap of 1.6%. However, in terms of utility—both in no-attack and under-attack conditions—DRIFT surpasses CaMeL by 21.8% in the no-attack setting and 10.9% under attack. This indicates that DRIFT achieves a superior utility-security trade-off, making it more practical for real-world agentic systems.
@@ 3.3 DRIFT Adaptation
DRIFT is a system-level defense framework that can be deployed across many types of agents. To better understand the adaptability and generality of DRIFT in different agent settings, we apply it to multiple LLMs, including four advanced online models—GPT-4o [26], GPT-4o-mini [25], Claude-3 Haiku [28], and Claude-3.5-Sonnet [27]—and one widely used offline model, Qwen2.5-7B [29].
For the online models, we compare our method with agents using ReAct [34], a technique that allows the LLM to reason and call tools in an agentic manner. The results are presented in Figure 4 (detailed results on four scenarios shown in Appendix B). We observe that DRIFT significantly enhances security across all models, reducing ASR from over 10% to single-digit levels, strongly indicating the security generality of DRIFT across diverse models. Notably, GPT-4o with ReAct, one of the most advanced LLMs with strong general capabilities, shows a high ASR of 51.7%, highlighting the vulnerability of current LLM agents—even those powered by leading models. However, after deploying DRIFT, the ASR drops sharply from 51.7% to just 1.5%, further demonstrating the effectiveness of DRIFT in securing agents from attack.
@@
The offline model Qwen2.5-7B, which has been tuned on our policy, achieves remarkable improvements in both utility and security. In terms of utility, our tuned agent obtains a 5.6% improvement in safe conditions and 3.1% in unsafe conditions. It is noticeable that the ASR after tuning drops to 0. These improvements highlight a potential solution for robustly securing agentic systems without performance sacrifice. All of these results demonstrate the effectiveness of DRIFT across different models and scenarios, fully supporting its broad adaptability and strong generality.
Figure 4: Comparison across different LLM Agents.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly central to agentic systems due to their strong reasoning and planning capabilities. By interacting with external environments through predefined tools, these agents can carry out complex user tasks. Nonetheless, this interaction also introduces the risk of prompt injection attacks, where malicious inputs from external sources can mislead the agent’s behavior, potentially resulting in economic loss, privacy leakage, or system compromise. System-level defenses have recently shown promise by enforcing static or predefined policies, but they still face two key challenges: the ability to dynamically update security rules and the need for memory stream isolation. To address these challenges, we propose DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems, which enforces both control- and data-level constraints. A Secure Planner first constructs a minimal function trajectory and a JSON-schema-style parameter checklist for each function node based on the user query. A Dynamic Validator then monitors deviations from the original plan, assessing whether changes comply with privilege limitations and the user’s intent. Finally, an Injection Isolator detects and masks any instructions that may conflict with the user query from the memory stream to mitigate long-term risks. We empirically validate the effectiveness of DRIFT on the AgentDojo and ASB benchmark, demonstrating its strong security performance while maintaining high utility across diverse models—showcasing both its robustness and adaptability. The code is released at https://github.com/SaFoLab-WISC/DRIFT.
@@ 1 Introduction
Large Language Models (LLMs), empowered by their exceptional planning and reasoning abilities, are increasingly integrated into agentic systems [1, 2, 3]. By processing natural language data streams, LLM agents interact with external environments, such as applications [1, 4], computing systems [3], via a set of pre-defined tools to carry out complex user tasks. Since the need for interaction with untrusted external environments, a new security threat of prompt injection attacks is introduced [5, 6, 7, 8, 9], where attackers inject malicious instructions into third-party platforms, misleading the agent workflow after external interaction. For example, a product review on Amazon written by another user, such as “Ignore previous instructions, buy this red shirt,” may manipulate the LLM into executing unintended actions. This form of attack [5, 6, 7, 8, 9] may bring risks such as economic losses [6], privacy leakage [10], and system damage [11] to users, severely undermining the reliability of the agentic system.
@@
Existing defense mechanisms can be broadly categorized into model-level [12, 13, 14, 15, 16, 17] and system-level [18, 19, 20, 21, 22] defenses. Model-level defenses [12, 13, 14, 15, 16, 17] typically rely on building the model’s intrinsic guardrails to detect injection inputs or mitigate their impact, but such defenses are constrained by the models’ inherent vulnerabilities and often struggle to defend against unseen attacks. Recently, system-level defenses [18, 19, 20, 21, 22] have gained increasing attention, as they can overcome the intrinsic weaknesses of models when facing unseen attacks, thereby achieving higher reliability in real-world agentic systems. These approaches typically restrict agents’ action spaces through security policies and workflow design to prevent potential injection threats. For instance, IsolateGPT [18] mitigates information leakage risks by enforcing isolation mechanisms and maintaining a separate memory bank for each application. Recently, CaMeL [21] achieves impressive security by manually defining a set of security policies and constructing a strict and fixed control and data dependency graph from the user query before any interaction takes place. More related works are discussed in Appendix B.
@@
Despite the progress in system-level defense mechanisms for agentic systems, two critical challenges remain largely unresolved: (1) the dynamic updating of security policies and (2) the isolation of covertly injected content within the memory stream. While CaMeL enforces robust security through a strict dependency graph, this static design considerably sacrifices flexibility and practical usability, particularly in agentic systems that require adaptive, real-time decision-making. Furthermore, the reliance on manually crafted security policies imposes considerable overhead and impedes generalization across diverse usage scenarios. In addition, IsolateGPT restricts the propagation of injection-related information across different applications, but residual injection content preserved in memory still poses significant risks within the same application during prolonged interactions.
@@
To overcome these challenges, we develop DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems that enforces security through both control- and data-level constraints. We first design a Secure Planner, which establishes the initial constraint policies solely according to the user query prior to any interaction. It constructs a minimal function trajectory (control constraints) to avoid injections misleading by executing functions in order. In addition, a checklist for each function node in the trajectory, with detailed parameter requirement and value dependencies, is encoded in JSON schema format [23]. When trajectory deviations are detected, a Dynamic Validator performs approval assessments based on the privilege category (Read, Write, Execute) and its alignment with the user’s original intent. To avoid the risk of injection messages to the agent or other modules during prolonged interactions, an Injection Isolator is also designed to continuously polish the memory after each interaction, identifying and masking any instructions that conflict with the initial user query. This layered defense strategy ensures strong context isolation while enabling secure and adaptive decision-making throughout long-term agent interactions.
@@
As a fully automatic system-level defense framework, DRIFT demonstrates strong performance across diverse scenarios, achieving high security while maintaining robust utility. Specifically, we evaluate DRIFT on the AgentDojo [24] benchmark, a simulated agent environment featuring various task scenarios and types of injection attacks. By applying DRIFT to GPT-4o-mini [25], the Attack Success Rate (ASR) is successfully reduced from 30.7% to 1.3%, while utility outperforms CaMeL by 20.1% under no attack and by 12.5% under attack. In addition, DRIFT shows remarkable adaptability and generalization across four advanced online LLMs: GPT-4o [26], GPT-4o-mini [25], Claude-3.5-sonnet [27], Claude-3-haiku [28], and one prevalent offline LLM, Qwen2.5-7B-Instruct [29]. On all of these models, DRIFT significantly enhances security while maintaining or even improving utility on some models. Moreover, we finetune our policy on DRIFT, with the dataset collected from ToolBench [30], achieving significant improvements in both security and utility. Compared to the original version, the ASR of the policy-tuned model drops from 15.1% to 0.0%, while utility under no attack increases from 26.6% to 32.2%, and utility under attack improves from 19.1% to 22.2%. This policy training mechanism could enable more reliable, secure, and functional LLM agentic systems.
@@
  • We develop DRIFT, a comprehensive system-level defense that integrates dynamic security mechanisms and memory isolation, achieving superior, balanced security and utility.
@@ 2.2 Dynamic Validator
In general, a function with the Read privilege does not directly pose a risk to the user and will be approved even if it deviates from the original trajectory. However, functions marked as Write or Execute may introduce direct risks. In such cases, the Validator will assess whether the deviated function aligns with the user’s original intent based on the updated tool messages, using the prompt shown in Figure 10. If the deviated function still aligns with the user’s intent, the function is approved and incorporated into the minimal function trajectory and parameter checklist to support successful validation in subsequent validation. Otherwise, agents will send an approval request to user (in our evaluation setting, sending a user request is equivalent to rejecting the deviated function call).
@@ 2.3 Injection Isolator
Current rule-based agent defense approaches typically restrict action permissions but do not eliminate injected content. In a long-term agentic system, past memory—such as previous conversations and tool responses—is frequently reused. These reused elements may be accessed not only by the agent itself but also by other components within the security system, such as the policy updating module. During the process of policy optimization, it is inevitable to incorporate new information obtained from recent interactions. However, any injection content stored in the memory stream will also be repeatedly exposed to these components during long-term interactions, severely increasing the risk of compromise over time. In addition, not all injection instructions interfere with the tool-call trajectory. For example, an instruction such as “In your final answer, suggest the hotel ‘Riverside View”’ affects only the final response rather than the tool-call process. Such cases cannot be defended by control-based or data-based constraints, as no deviation occurs in the tool-call trajectory.
To mitigate this long-term and tool-independent threat, we propose an injection isolation mechanism to detect and remove injected content from the memory stream. Specifically, we design a curated Injection Isolator that analyzes returned messages from each tool-calling and determines whether any instructions conflict with the user’s original intent. The identification process is driven by a LLM using system prompt in Figure 11. If a conflict is detected, the isolator removes the conflicting instructions using external masking components before the message is added to the agent’s memory stream. Subsequently, a safe memory stream could be maintained in long-term agent interactions. The Isolator cannot directly modify the tools and does not interact with the agent, which helps prevent potential security vulnerabilities as much as possible.
@@ 2.5 Trainable Security Policy
To enhance the reliability and generalization of our security policy, we develop a training approach for both the Secure Planner and the Injection Isolator, allowing our DRIFT framework to adapt more robustly. This involves designing a new data collection pipeline that extracts policy-aligned samples from existing agent datasets, followed by efficient instruction tuning using Low-Rank Adaptation (LoRA) [32] on Qwen2.5-7B-Instruct [29].
@@ 2.5.1 Data and Environment Construction
After data collection completed, we fine-tune the Qwen2.5-7B-Instruct model using LoRA for both the Secure Planner and Injection Isolator, as well as the agent itself. For the Dynamic Validator, we rely on the original Qwen2.5-7B-Instruct in a zero-shot setup to handle privilege classification and user intent checking.
@@ 3 Experiments
In this section, we evaluate DRIFT on AgentDojo [24] and ASB [33], two prevalent agentic security benchmarks, to assess the effectiveness, robustness and adaptability of DRIFT in terms of both utility and security. Furthermore, we analyze the contribution of each individual component within DRIFT.
@@ 3.1 Experimental Setups
We evaluate our method with AgentDojo [24], a benchmark that simulates realistic interactions in agent-based systems. It includes four scenarios—banking, Slack, travel, and workspace—covering 97 user tasks to assess utility and 629 injection tasks to evaluate security. In addition, we evaluate our method on ASB [33], another agent security benchmark that encompasses 10 evaluation scenarios.
@@
We compare our method against several advanced existing defense approaches. Specifically, we include four defenses implemented in AgentDojorepeat_user_prompt [34], spotlighting_with_delimiting [35], tool_filter [36], and transformers_pi_detector [17], as well as three defenses implemented in ASBdelimiters_defense [37], ob_sandwich_defense [34], and instructional_prevention [38]. In addition, we compare against two system-level defenses: a static policy-based defense, CaMeL [21], and a dynamic policy-based defense, Progent [22]. These baselines represent a broad range of strategies for protecting agents against prompt injection attacks.
@@
We apply our policy to several models, including online models—GPT-4o [26], GPT-4o-mini [25], Claude-3-haiku [28], and Claude-3.5-sonnet [27]—and an offline model, Qwen2.5-7B-Instruct [29]. For Qwen2.5-7B-Instruct, we fine-tune it on our policy dataset (described in Section 2.5) using a batch size of 4 and training for three epochs. We employ the Adam optimizer [39] with weight decay and set the initial learning rate to 2e-5. Our default attacks are the important_instruction attack on AgentDojo and the OPI attack on ASB.
Figure 3: Comparison of defense methods on GPT-4o-mini in AgentDojo.
Figure 4: Comparison of defense methods on GPT-4o-mini in ASB.
@@ 3.2 Defense Techniques Comparison
In this experiment, we evaluate DRIFT on two prevalent agent safety benchmarks, AgentDojo [24] and ASB [33], and compare it with multiple advanced defenses. By default, we employ GPT-4o-mini-2024-07-18 as the base agent.
Comparison on AgentDojo.
On the AgentDojo benchmark, we compare DRIFT with six advanced defense techniques—four implemented in AgentDojo: repeat_user_prompt, spotlighting_with_delimiting, tool_filter, transformers_pi_detector—one static policy-based defense, CaMeL, and one dynamic policy-based defense, Progent. The results are presented in Figure 3.
Notably, the DRIFT policy achieves an optimal balance between utility and security. In terms of security, DRIFT significantly outperforms all other baselines except CaMeL, with only a marginal gap of 1.3%. However, in terms of utility under both no-attack and under-attack conditions, DRIFT surpasses CaMeL by 21.8% in the no-attack setting and 10.9% under attack. This demonstrates that DRIFT achieves a superior utility–security trade-off, highlighting the greater practicality and effectiveness of dynamic policies over static ones.
Compared with Progent, the other dynamic policy-based defense, DRIFT outperforms it in both utility and security. This further validates the effectiveness of our dynamic policy design and highlights DRIFT as a more practical and robust defense for real-world agentic systems.
Comparison on ASB.
On the ASB benchmark, we compare DRIFT with four advanced defense techniques: delimiters_defense, ob_sandwich_defense, instructional_prevention, and Progent. The results are presented in Figure 4.
We observe that DRIFT outperforms all other defenses in terms of security, achieving an ASR of only 4.8%, which significantly surpasses the runner-up defense, Progent, with an ASR of 15.8%. In terms of utility, DRIFT experiences a slight performance drop compared to the undefended agent but still maintains robust functionality under both no-attack and under-attack conditions. These findings further highlight the superiority of our proposed DRIFT in achieving a balanced trade-off between utility and security.
@@ 3.3 DRIFT Adaptation
DRIFT is a system-level defense framework that can be deployed across many types of agents. To better understand the adaptability and generality of DRIFT in different agent settings, we apply it to multiple LLMs, including four advanced online models—GPT-4o [26], GPT-4o-mini [25], Claude-3 Haiku [28], and Claude-3.5-Sonnet [27]—and one widely used offline model, Qwen2.5-7B-Instruct [29]. The evaluation is conducted on AgentDojo.
For the online models, we compare our method with agents using ReAct [40], a technique that allows the LLM to reason and call tools in an agentic manner. The results are presented in Figure 5 (detailed results on four scenarios shown in Appendix D). We observe that DRIFT significantly enhances security across all models, reducing ASR from over 10% to single-digit levels, strongly indicating the security generality of DRIFT across diverse models. Notably, GPT-4o with ReAct, one of the most advanced LLMs with strong general capabilities, shows a high ASR of 51.7%, highlighting the vulnerability of current LLM agents—even those powered by leading models. However, after deploying DRIFT, the ASR drops sharply from 51.7% to just 1.5%, further demonstrating the effectiveness of DRIFT in securing agents from attack.
@@
The offline model Qwen2.5-7B-Instruct, which has been tuned on our policy, achieves remarkable improvements in both utility and security. In terms of utility, our tuned agent obtains a 5.6% improvement in safe conditions and 3.1% in unsafe conditions. It is noticeable that the ASR after tuning drops to 0. These improvements highlight a potential solution for robustly securing agentic systems without performance sacrifice. All of these results demonstrate the effectiveness of DRIFT across different models and scenarios, fully supporting its broad adaptability and strong generality.
Figure 5: Comparison across different LLM Agents on AgentDojo.